### v2

@@ 1 Introduction
Figure 1: In GBP Learning, we design factor graphs whose structure mirrors common NN architectures, enabling distributed training and prediction with GBP. Learnable parameters are included as random variables (circles), as are inputs, outputs and activations. The parameters are shared over across all observations, where the other variables are copied once per observation. Factors (black squares) between layers constrain their representations to be locally consistent, while those attached to inputs and outputs encourage compatibility with observation. Non-linear inter-layer factors to enable soft-switching behaviour. This example architecture for image classification comprises convolutional, max pooling and dense projection layers. The same architecture could learn in an unsupervised way by removing the output observation factor.
@@
Concurrently, we reflect on how neural networks (NNs) are trained. Despite backpropagation rumelhart1985learning being largely successful for DL, training NNs spread over multiple processors is made difficult by backward locking: processors for earlier layers sit idle after their part in the forward pass, awaiting the backward error signal. This challenge will become more pertinent with the growth of i) larger models which must be distributed over many devices, ii) new hardware architectures whose cores have significant local memory (Graphcore; cerebras), and iii) parallel, distributed and heterogeneous embedded devices (Sutter:Jungle2011). We thus expect a growing need for more flexible training algorithms which admit efficient model-parallelism.
@@
Much recent work has shown that Gaussian BP (GBP) to be a robust and effective algorithm for distributed inference in factor graphs, even in the presence of non-linear and non-Gaussian factors (Davison:Ortiz:ARXIV2019; Ortiz:etal:CVPR2020; Murai:etal:ARXIV2022; Patwardhan:etal:ARXIV2022). It is thus our choice of inference engine here. By approximating the factors in our model as Gaussian, we can use of GBP for training (inference over parameters, given observations) and prediction (inference of e.g. outputs, given parameters and inputs). The generality of GBP provides flexibility as to which variables are observed, meaning training and prediction are essentially the same computation, and partial observations or missing labels do not require fundamentally different treatment. As BP is inherently local and stateful, our training can be distributed and asynchronous.
Our work shares similar goals to lucibello2022deep(lucibello2022deep), who train MLP-like factor graphs with GBP using analytically derived message updates. However, their approach lacks generality: the derived updates hold only for dense layers, and they focus on models with binary weights and sign activation functions. Here, in contrast, we present a general method to train deep models with GBP, requiring only that appropriate energy functions be specified for inter-layer factors, and naturally supporting continuous weights. Message updates need not be rederived for different architectures.
We call our approach GBP Learning. In short, we construct factor graph architectures inspired by those in DL (Section 3.1). Within these models, GBP inference admits flexible and distributed training and prediction (Section 3.2) and we can learn incrementally via Bayesian filtering over parameters (Section 3.4). Our experiments demonstrate the benefit of factor graphs with learnable parameters over hand-designed solvers in a video denoising task (Section 5.2). For image classification (Section 5.3), we evaluate our continual learning approach by single epoch training on MNIST. We achieve performance equivalent to an Adam-trained CNN with a replay buffer of $6\times 10^{3}$ examples, and show this performance is robust to asynchronous training. Last, we compare to lucibello2022deep(lucibello2022deep), outperforming them on both MNIST (by $0.8\%$) and CIFAR10 (by $11.8\%$).
@@
    A new class of deep networks which we represent as Gaussian factor graphs and train with GBP.
@@ 2.1 Factor graphs
Here $\mathbf{x}_{\phi_{j}}$ denotes the vector of all $V_{j}$ variables in the neighbourhood of $\phi_{j}$ and $Z$ is a normalising constant. A factor graph is bipartite: variables only connect to factors and vice versa. Each factor may encode an observation, or prior about one of many variables. Factor functions $\phi_{j}(\cdot)$ may be unnormalised distributions relating to their energy, $E_{j}(\cdot)$, as
@@ 2.2 Belief Propagation
Belief Propagation (pearl1988probabilistic) is a message passing algorithm which can perform inference on factor graphs via distributed, iterative computation. Each message is passed along the edge between one factor $\phi_{j}$ and one variable $x_{i}$. Messages travel in both directions, and we use the notations $m_{{\phi_{j}\rightarrow x_{i}}}$ and $m_{{x_{i}\rightarrow\phi_{j}}}$ for the two types of message. Each message is a probability distribution in the space of the variable involved. The following update rules are iterated until convergence:
@@
where $\mathrm{ne}\left(i\right)\setminus j$ is the indices of all nodes connected to $i$, except $j$. After convergence the posterior marginal of a variable is estimated by the product of its incoming messages:
@@
where $Z_{i}$ is straightforward to compute if the messages belong to a known parametric family. In this work, we assume the variable nodes are univariate, but BP can be extended for multivariate posterior inference by passing vector messages between sets of variables and factors. BP updates (4), (5), (6) have a number of interesting properties. First, the required computations rely only on the local state of the graph; no global context is necessary. Second, for exponential family distributions under natural parameterisation, taking the product of a set of messages is reduced to the addition of their parameters.
@@ 2.2.1 Gaussian Belief Propagation
where we have used $\Sigma_{\setminus i,\setminus i}^{\left(\phi_{j}+m\right)}=\left(\Lambda^{\left(\phi_{j}\right)}_{\setminus i,\setminus i}+\left(\mathbf{D}_{\phi_{j}}\right)_{\setminus i,\setminus i}\right)^{-1}$, $\eta^{\left(\phi_{j}+m\right)}:=\eta^{\left(\phi_{j}\right)}+\mathbf{e}_{\phi_{j}}$, $\mathbf{D}_{\phi_{j}}$ for the matrix of precision messages coming into $\phi_{j}$, and $\mathbf{e}_{\phi_{j}}$ for the vector of incoming information messages. Subscript ${}_{\setminus r}$ indicates all elements except that for variable $r$. Note that $\mathbf{D}_{\phi_{j}}$ is diagonal for graphs with scalar variable nodes.
As in the general case, GBP is not guaranteed to converge in graphs with cycles. However, if it does converge in linear-Gaussian models, it is guaranteed to converge to the correct posterior means (weiss1999correctness). In practice, it has been empirically shown that it converges to correct means in in many types of factor graph. We refer the reader to ortiz2021visual for an in-depth introduction to GBP.
@@ 2.3 Non-linear Factors
GBP supports the use of factors which include non-linear transformations of their connected variables, and this is crucial in our proposed use here for representation learning. Every time a message from a non-linear factor is computed, the factor is linearised around the current estimates of the variables, $\mathbf{X}_{0}$, i.e. for each factor $\mathbf{h}\left(\mathbf{x}_{\phi_{j}}\right)\approx\mathbf{h}\left(\mathbf{x}_{\phi_{j},0}\right)+\mathbf{J}_{j}^{{}^{\top}}\left(\mathbf{x}_{\phi_{j}}-\mathbf{x}_{\phi_{j},0}\right)$ where $\mathbf{J}_{j}:=\left.\partial\mathbf{h}\left(\mathbf{x}_{\phi_{j}}\right)/\partial\mathbf{x}_{\phi_{j}}\right|_{\mathbf{x}_{\phi_{j},0}}$. Linearisation can be implemented using the following forms of factor precision and information in (8):
$\Lambda^{(\phi_{j})}$ $\approx\mathbf{J}_{j}^{{}^{\top}}\Lambda_{\mathbf{y}_{j}}\mathbf{J}_{j}$ (9)
$\eta^{(\phi_{j})}$ $\approx\mathbf{J}_{j}^{{}^{\top}}\Lambda_{\mathbf{y}_{j}}\left(\mathbf{J}_{j}^{{}^{\top}}\mathbf{x}_{\phi_{j},0}+\mathbf{y}_{j}-\mathbf{h}\left(\mathbf{x}_{\phi_{j},0}\right)\right).$
See Section 3.3 of Davison:Ortiz:ARXIV2019 for details.
@@ 3 GBP Learning
Our aim is to produce factor graphs which have similar architectural inductive biases to NNs, that can be overparameterised in a similar way, but can be trained with GBP. We will now describe the key factors in our model and our efficient GBP routine for training and prediction. The factor energies of form (3) contain $\mathbf{h}\left(\cdot\right)$ and $\mathbf{y}$ which, along with $\Lambda_{\mathbf{y}}$, are sufficient to deduce the linearised factor (9). This, in turn, defines the GBP message updates (7), (8). It is thus sufficient to describe our model in terms of factor energies.
@@ 3.1 Deep Factor Graphs
$r\left(\mathbf{A},\mathbf{B},k\right)$ $:=g\left(\mathrm{vec}\left(A\right)\cdot\mathrm{vec}\left(B\right)+k\right)~{}.$ (12)
$g(\cdot)$ is an elementwise nonlinear activation function, and we have removed the functional dependence of $E_{\mathrm{conv}}^{(a,b,c)}$ on the connected variables for brevity. The total energy of the layer is found by summing $E_{\mathrm{conv}}^{(a,b,c)}$ over spatial locations $(a,b)$ and output channels $c$.
We can similarly define a transposed convolution layer. In this case, each filter is weighted by the output variable of its corresponding channel, and the weighted sum reconstructs the inputs. Note that for a stride smaller than the kernel size, each input will belong to multiple receptive fields and they are summed over to give its reconstruction. For example, for a stride of one and neglecting edge effects, these contributions can be combined according to:
$E_{\mathrm{convT}}^{(a,b,c)}$ $=\frac{\left(x_{l-1}^{(a,b,c)}-r\left(\mathbf{X}_{l}^{(a,b)},\mathbf{\theta}_{l}^{(c)},b_{l}^{(c)}\right)\right)^{2}}{2\sigma_{l}^{2}}~{}.$
Note that in this case the parameters have dimensions $\mathbf{\theta}_{l}^{(c)}\in\mathbb{R}^{K_{l}\times K_{l}\times C_{l}}$ and $b_{l}^{(c)}\in\mathbb{R}^{C_{l-1}}$.
For image classification, we use a dense projection head which translates the activation variables for the preceding layer $\mathbf{X}_{L-1}\in\mathbb{R}^{H_{L-1}\times W_{L-1}\times C_{L-1}}$ to a vector $\mathbf{x}_{L}\in\mathbb{R}^{C_{L}}$, where $C_{L}$ is the number of image classes. The dense factor energy is:
@@
where the activation function $g(\cdot)$ is included for generality, but set to identity for image classification.
@@
These “layer” abstractions can be composed to produce deep models with similar design freedom to DL. Further, our models may be overparameterised by introducing large numbers of learnable weights and we believe that the inclusion of non-linear activation functions $\mathbf{g}(\cdot)$ in our factor graph can aid representation learning as in DL. In particular, we note that for non-linear factors, the factor Jacobian is a function of the current variable values $\mathbf{J}_{j}=\mathbf{J}_{j}\left(\mathbf{x}_{\phi_{j},0}\right)$, which will cause the strength of the factors (9) to vary depending on the input data. We expect this to produce a similar “soft switching” of connections as observed in non-linear NNs.
@@ 3.2 Learning and Predicting with GBP Inference
We have described the components of deep factor graph models, in which inputs, outputs, activations and parameters are random variables. As we have designed these models with (non-linear) Gaussian factors, we can do inference with GBP via iterative linearisation (Section 2.3). We apply this inference engine to estimate posteriors over all latent variables. Note that there is no fundamental difference between training and prediction — only a difference in which variables are observed. In training, we infer a posterior over parameters given observed inputs and, where supervision is available, outputs. To then make predictions on new examples, we run GBP to predict the unobserved inputs/outputs given the observed inputs/outputs and parameters.
Our message schedule proceeds as follows. For a given batch, we initialise the graph and update messages by sweeping forward and backward through the layers: first nearest the input observations, progressing to the deepest layer and back again. We repeat this for a specified number of iterations. Within each layer, we compute all factor to variable updates in parallel, and likewise for the variable to factor messages of each variable type (inputs, outputs, parameters as applicable). In addition, we experimentally show our approach works well with layerwise asynchronous message schedules (Section 5.3.1). While we have demonstrated these schedules work, they are likely suboptimal and we leave exploration for future work. We find that applying damping (murphy2013loopy) and dropout to the factor to variable messages is sufficient for stable GBP.

### v3

@@ Abstract
Belief Propagation, Bayesian Deep Learning, Energy Based Models, Local Learning
@@ 1 Introduction
Figure 1: In GBP Learning, we design factor graphs whose structure mirrors common NN architectures, enabling distributed training and prediction with GBP. Learnable parameters are included as random variables (circles), as are inputs, outputs and activations. The parameters are shared over across all observations, where the other variables are copied once per observation. Factors (black squares) between layers constrain their representations to be locally consistent, while those attached to inputs and outputs encourage compatibility with observation. The inter-layer factors are non-linear to enable soft-switching behaviour. This example architecture for image classification comprises convolutional, max pooling and dense projection layers. The same architecture could be trained without supervision by removing the output observation factor.
@@
Concurrently, we reflect on how neural networks (NNs) are trained. Despite backpropagation (rumelhart1985learning) being largely successful for DL, training NNs spread over multiple processors is made difficult by backward locking: processors for earlier layers sit idle after their part in the forward pass, awaiting the backward error signal. This challenge will become more pertinent with the growth of i) larger models which must be distributed over many devices, ii) new hardware architectures whose cores have significant local memory (Graphcore; cerebras), and iii) parallel, distributed and heterogeneous embedded devices (Sutter:Jungle2011). We thus expect a growing need for more flexible training algorithms which admit efficient model-parallelism.
@@
Much recent work has shown that Gaussian BP (GBP) to be a robust and effective algorithm for distributed inference in factor graphs, even in the presence of non-linear and non-Gaussian factors (Davison:Ortiz:ARXIV2019; Ortiz:etal:CVPR2020; Murai:etal:ARXIV2022; Patwardhan:etal:ARXIV2022). It is thus our choice of inference engine here. By approximating the factors in our model as Gaussian, we can use GBP for training (inference over parameters, given observations) and prediction (inference of e.g. outputs, given parameters and inputs). The generality of GBP provides flexibility as to which variables are observed, meaning training and prediction are essentially the same computation, and partial observations or missing labels do not require fundamentally different treatment. As BP is inherently local and stateful, our training can be distributed and asynchronous.
Our work shares similar goals to lucibello2022deep(lucibello2022deep), who train MLP-like factor graphs using GBP with analytically derived message updates. For computational efficiency, they focus on architectures with binary weights and sign activation functions. In contrast, our approach enables training with GBP on arbitrary architectures, without rederivation of the message updates; and we focus on models with continuous weights in natural analogue to NNs. We demonstrate this generality, training convolutional architectures with our approach and showing we can outperform lucibello2022deep(lucibello2022deep) on image classification tasks.
We call our approach GBP Learning. Our models (Section 3.1) are factor graphs with architectures inspired by those in DL, and factors between layers to enforce multi-layer consistency as used in the predictive coding literature (millidge2022predictive). Within these models, GBP admits flexible and distributed training and prediction (Section 3.2), and learning can be done incrementally via Bayesian filtering over parameters (Section 3.4). Our experiments demonstrate the benefit of factor graphs with learnable parameters over hand-designed solvers in a video denoising task (Section 5.2). For image classification (Section 5.3), we evaluate our continual learning approach by single epoch training on MNIST. We achieve performance equivalent to an Adam-trained CNN with a replay buffer of $6\times 10^{3}$ examples, and show this performance is robust to asynchronous training. Last, we compare to lucibello2022deep(lucibello2022deep), outperforming them on both MNIST (by $0.8\%$) and CIFAR10 (by $11.8\%$).
@@
    An approach to train deep factor graphs with GBP, which can be applied to any architecture and supports incremental learning.
@@ 2.1 Factor graphs
Here $\mathbf{x}_{\phi_{j}}$ denotes the vector of all $V_{j}$ variables in the neighbourhood of $\phi_{j}$ and $Z$ is a normalising constant. A factor graph is bipartite: variables only connect to factors and vice versa. Each factor may encode an observation, or prior on one or many variables. Factor functions $\phi_{j}(\cdot)$ may be unnormalised distributions relating to their energy, $E_{j}(\cdot)$, as
@@ 2.2 Belief Propagation
Belief Propagation (pearl1988probabilistic) is a message passing algorithm to perform inference on factor graphs via distributed, iterative computation. Each message is passed along the edge between one factor $\phi_{j}$ and one variable $x_{i}$. Messages travel in both directions, and we use the notations $m_{{\phi_{j}\rightarrow x_{i}}}$ and $m_{{x_{i}\rightarrow\phi_{j}}}$ for the two message types. Each message is a probability distribution in the space of the variable involved. The following update rules are iterated until convergence:
@@
where $\mathrm{ne}\left(k\right)\setminus l$ denotes the indices for variables connected to factor $\phi_{k}$, except $l$. After convergence the posterior marginal of a variable is estimated by the product of its incoming messages:
@@
where $Z_{i}$ is straightforward to compute if the messages belong to a known parametric family. In this work, we assume the variable nodes are univariate, but BP can be extended for multivariate posterior inference by passing vector messages between sets of variables and factors.
BP updates (4), (5), (6) have a number of interesting properties. First, the required computations rely only on the local state of the graph; no global context is necessary. Second, for exponential family distributions under natural parameterisation, taking the product of a set of messages is reduced to the addition of their parameters.
@@ 2.2.1 Gaussian Belief Propagation
where we have used $\Sigma_{\setminus i,\setminus i}^{\left(\phi_{j}+m\right)}=\left(\Lambda^{\left(\phi_{j}\right)}_{\setminus i,\setminus i}+\left(\mathbf{D}_{\phi_{j}}\right)_{\setminus i,\setminus i}\right)^{-1}$, $\eta^{\left(\phi_{j}+m\right)}:=\eta^{\left(\phi_{j}\right)}+\mathbf{e}_{\phi_{j}}$, $\mathbf{D}_{\phi_{j}}$ for the matrix of precision messages coming into $\phi_{j}$, and $\mathbf{e}_{\phi_{j}}$ for the vector of incoming information messages. Subscript ∖r indicates all elements except that for variable $r$. Note that $\mathbf{D}_{\phi_{j}}$ is diagonal for graphs with scalar variable nodes.
Convergence and correctness of GBP and similar algorithms has been studied extensively (e.g. malioutov2006walk; moallemi2009convergence; moallemi2010convergence). As for general models, BP is not guaranteed to converge in Gaussian models with cycles. However, if it does converge in linear-Gaussian models, it is guaranteed to converge to the correct posterior means (weiss1999correctness). Further, many iterative pre-conditioning schemes have been proposed which guarantee convergence in the linear-Gaussian case (johnson2009fixing; ruozzi2013message). The guarantees do not hold for models with non-linear factors. We refer the reader to ortiz2021visual(ortiz2021visual) for an intuitive introduction to GBP.
@@ 2.3 Non-linear Factors
GBP supports the use of factors which include non-linear transformations of their connected variables, and this is crucial in our proposed use here for representation learning. Every time a message from a non-linear factor is computed, the factor is linearised around the current variable estimates, $\mathbf{X}_{0}$, i.e. for each factor $\mathbf{h}\left(\mathbf{x}_{\phi_{j}}\right)\approx\mathbf{h}\left(\mathbf{x}_{\phi_{j},0}\right)+\mathbf{J}_{j}^{{}^{\top}}\left(\mathbf{x}_{\phi_{j}}-\mathbf{x}_{\phi_{j},0}\right)$ where $\mathbf{J}_{j}:=\left.\partial\mathbf{h}\left(\mathbf{x}_{\phi_{j}}\right)/\partial\mathbf{x}_{\phi_{j}}\right|_{\mathbf{x}_{\phi_{j},0}}$. The resulting approximated factor has a quadratic (Gaussian) energy, and the following factor precision and information can be derived
$\Lambda^{(\phi_{j})}$ $\approx\mathbf{J}_{j}^{{}^{\top}}\Lambda_{\mathbf{y}_{j}}\mathbf{J}_{j}~{};$ (9)
$\eta^{(\phi_{j})}$ $\approx\mathbf{J}_{j}^{{}^{\top}}\Lambda_{\mathbf{y}_{j}}\left(\mathbf{J}_{j}^{{}^{\top}}\mathbf{x}_{\phi_{j},0}+\mathbf{y}_{j}-\mathbf{h}\left(\mathbf{x}_{\phi_{j},0}\right)\right)~{}.$
(9) can be substituted into the message update rules (8), enabling GBP inference in the linear-approximated model.
To do approximate inference in the original model, one can alternate between i) setting the linearisation point $\mathbf{X}_{0}$ to be the MAP estimate of the variables given current messages, and ii) updating messages given the current linearisation point. We emphasise that only non-linear factors need to be iteratively approximated in this way, and linear-Gaussian factors remain static throughout inference. We refer the reader to Section 3.3 of (Davison:Ortiz:ARXIV2019) for further details on GBP in non-linear models.
@@ 3 GBP Learning
Our aim is to produce factor graphs which have similar architectural inductive biases to NNs, that can be overparameterised in a similar way, but can be trained with GBP. We will now describe the key factors in our model and our efficient GBP routine for training and prediction. The factor energies of form (3) contain $\mathbf{h}\left(\cdot\right)$ and $\mathbf{y}$ which, along with $\Lambda_{\mathbf{y}}$, are sufficient to deduce the linearised factor (9). This, in turn, defines the GBP message updates (7), (8). It is thus sufficient to describe our model in terms of factor energies. We emphasise that our model parameters are included as random variables in the factor graph and inferred with GBP.
@@ 3.1 Deep Factor Graphs
$r\left(\mathbf{A},\mathbf{B},s\right)$ $:=g\left(\mathrm{vec}\left(A\right)\cdot\mathrm{vec}\left(B\right)+s\right)~{}.$ (12)
$g(\cdot)$ is an elementwise non-linear activation function, and we have removed the functional dependence of $E_{\mathrm{conv}}^{(a,b,c)}$ on the connected variables for brevity. The total energy of the layer is found by summing $E_{\mathrm{conv}}^{(a,b,c)}$ over spatial locations $(a,b)$ and output channels $c$.
We can similarly define a transposed convolution layer. In this case, each filter is weighted by the output variable of its corresponding channel, and the weighted sum reconstructs the inputs. Note that for a stride smaller than the kernel size, each input will belong to multiple receptive fields which are summed over to give its reconstruction. For example, for a stride of one and neglecting edge effects, these contributions can be combined according to:
$E_{\mathrm{convT}}^{(a,b,c)}$ $=\frac{\left(x_{l-1}^{(a,b,c)}-r\left(\mathbf{X}_{l}^{(a,b)},\mathbf{\theta}_{l}^{(c)},b_{l}^{(c)}\right)\right)^{2}}{2\sigma_{l}^{2}}~{}.$ (13)
In this case, the parameters have dimension $\mathbf{\theta}_{l}^{(c)}\in\mathbb{R}^{K_{l}\times K_{l}\times C_{l}}$ and $b_{l}^{(c)}\in\mathbb{R}$.
For image classification, we use a dense projection layer which translates the activations in the preceding layer $\mathbf{X}_{L-1}\in\mathbb{R}^{H_{L-1}\times W_{L-1}\times C_{L-1}}$ to a vector $\mathbf{x}_{L}\in\mathbb{R}^{C_{L}}$, where $C_{L}$ is the number of classes. The dense factor energy is:
@@
where the activation function $g(\cdot)$ is included for generality, but usually chosen to be identity for a last-layer classifier.
@@
These “layer” abstractions can be composed to produce deep models with similar design freedom to DL. Further, our models may be overparameterised by introducing large numbers of learnable weights and we believe that the inclusion of non-linear activation functions $g(\cdot)$ in our factor graph can aid representation learning as in DL. In particular, we note that for non-linear factors, the factor Jacobian is a function of the current variable estimates $\mathbf{J}_{j}=\mathbf{J}_{j}\left(\mathbf{x}_{\phi_{j},0}\right)$, which will cause the strength of the factors (9) to vary depending on the input data. We expect this to produce a similar “soft switching” of connections as observed in non-linear NNs.
@@ 3.2 Learning and Predicting with GBP Inference
We have described the components of deep factor graph models, in which inputs, outputs, activations and parameters are random variables. As these models include non-linear factors such as (11), (13) and (14), we use the iterative linearisation scheme described in Section 2.3 to do approximate inference with GBP. We apply this inference engine to estimate posteriors over all latent variables. Note that there is no fundamental difference between training and prediction — only a difference in which variables are observed. In training, we infer a posterior over parameters given observed inputs and, where supervision is available, outputs. To then make predictions on new examples, we run GBP to predict the unobserved inputs/outputs given the observed inputs/outputs and parameters.
Our message schedule proceeds as follows unless stated otherwise. For a given batch, we initialise the graph and update messages by sweeping forward and backward through the layers: first nearest the input observations, progressing to the deepest layer and back again. We repeat this for a specified number of iterations. Within each layer, we compute all factor to variable updates in parallel, and likewise for the variable to factor messages of each variable type (inputs, outputs, parameters as applicable). In addition, we experimentally show our approach works well with layerwise asynchronous message schedules (Section 5.3.1). While we have demonstrated these schedules work, they are likely suboptimal and we leave exploration for future work. We find that applying damping (murphy2013loopy) and dropout to the factor to variable messages is sufficient for stable GBP.